40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 16 04:46:37 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
    Reply  Thu May 16 23:48:39 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
       Reply  Mon May 20 17:49:36 2013, Jenne, Update, LSC, PRMI sensing matrix - not high quality data PRMIsensMatQuickAnalysis.m.gz
          Reply  Mon May 20 18:47:47 2013, Koji, Update, LSC, PRMI sensing matrix - not high quality data 
             Reply  Tue May 21 14:48:08 2013, Jenne, Update, LSC, PRMI sensing matrix - not high quality data 
       Reply  Mon May 20 18:50:22 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
          Reply  Tue May 21 18:22:18 2013, Jenne, Update, LSC, Sensing matrix scripts modified to include actuator calibration 
             Reply  Wed May 22 00:08:19 2013, Koji, Update, LSC, Sensing matrix scripts modified to include actuator calibration 
                Reply  Wed May 22 18:24:19 2013, Jenne, Update, SUS, Violin mode survey 
                Reply  Wed May 22 20:50:26 2013, Jenne, Update, LSC, Sensing matrix scripts don't calculate correctly 
                   Reply  Thu May 23 14:45:08 2013, Jenne, Update, LSC, Sensing matrix scripts calculations make more sense now SensMat_JenneMeas_23May2013.pngSensMat_KojiMeas_23May2013.png
                      Reply  Thu May 23 21:45:51 2013, Jenne, Update, LSC, Sensing matrix scripts now check for lockloss 
                Reply  Thu May 23 19:09:15 2013, Jenne, Update, LSC, Sensing matrix scripts modified to include actuator calibration SensMat_21May2013.png
          Reply  Wed May 22 18:07:36 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
             Reply  Thu May 23 00:49:13 2013, Jenne, Update, LSC, LSC filters loaded 
Message ID: 8611     Entry time: Wed May 22 00:08:19 2013     In reply to: 8609     Reply to this: 8620   8621   8632
Author: Koji 
Type: Update 
Category: LSC 
Subject: Sensing matrix scripts modified to include actuator calibration 

It was too embarassing to see that the actuation frequency was set at the violin mode frequency in order to avoid designing a new filter!?

I ran Jenne's sensing matrix code and the immitated the same result by manual measurement with DTT.
I noticed that the PRM excitation was not transmitted to the mirror. I tracked down the cause and found that
Jenne is using 628Hz which is the notch frequency of the viloing filter.

There is no way we can measure the precise calibration of the error signal exactly at the violin mode frequency.

Nevertheless I waited for the ringdown of the violin mode to the floor level and ran the code again WITH the violin mode filter OFF
at PRM SUS.

The result was stored in the data file

sensematPRM_2013-05-22.12615.dat

The code spit the message at the end

Sensing Matrix, magnitude only, units = cts/meter
 
            MICH          PRCL         
AS55        5.304E+08     1.716E+09     
REFL11      1.732E+13     2.151E+11     
REFL33      1.616E+11     5.384E+09     
REFL55      1.681E+11     6.950E+09

Now I replicated the same measurement with DTT.

MICH or PRCL were excited with the lockin. In order to aviod the violin mode, I shifted the excitation freq by 1Hz. (i.e. 629.125Hz)

The peaks in REFL33I/Q and RFL55I/Q were observed with PSD and TF. The spectrum was measured with the FLATTOP window with the line resolution of 0.1Hz
DTT suggested that this corresponds to the BW of 0.471271Hz if I correctly understood what DTT plot said. We need this information to convert cnt/rtHz to cnt_pk
if we need. For the TF measurements, I needed to find the excitatin monitor but I could not. Therefore, I set the offset of LSC-LOCKIN1_SIG to be 1000,
so that C1:LSC-LOCKIN1_I_IN1 produce the same signal as the excitation.

Note that during the measurement, 628Hz nothces in the LSC servos were on. I confirmed that this provides the reduction of the feedback by a factor of 76.
As the original openloop gain at 629Hz is lower than the unity more than a factor of 2, this was sufficient attenuation to measure the optical gain with the systematic error of less than a %.

MICH excitation (ITMX -1, ITMY +1)
        PSD (cnt/rtHz)   TF Mag    Phase

REFL33I 0.098590         9.5691e-5 74.4344
REFL33Q 0.019294         1.8665e-5 71.1204
REFL55I 0.016123         1.3890e-5 77.3132
REFL55Q 0.157522         1.5285e-4 91.5594

PRCL excitation (PRM +1)
        PSD (cnt/rtHz)   TF Mag    Phase

REFL33I 15.7565          1.5298e-2 -109.727
REFL33Q  0.171648        1.6310e-4 -141.73
REFL55I 16.2834          1.5809e-2 -109.672
REFL55Q  0.634096        6.1012e-4 -143.169

These measurements are saved in the XML files (for DTT) in
/cvs/cds/caltech/users/koji/130521/
as
130521_MICH_EXC.xml and 130521_PRCL_EXC.xml

As the actuator of the PRM/ITMX/ITMY are {19.6, 4.70, 4.66}/f^2 nm/cnt, the optical gains were calculated from the TF measurements.

MICH excitation (ITMX -1, ITMY +1)
        OPTICAL GAIN
(cnt/m)
REFL33I 4.0e9
REFL33Q 7.9e8
REFL55I 5.9e8
REFL55Q 6.5e9

PRCL excitation (PRM +1)
        OPTICAL GAIN

REFL33I 3.1e11
REFL33Q 3.3e9
REFL55I 3.2e11
REFL55Q 1.2e10

These should be compared with the measurement by the script and we get more information from the script (like AS55, REFL11)

ELOG V3.1.3-